OpenStudy (anonymous):

find the domain of y=x+2/(x-3)(x+1) We are not as to the format of the answer

OpenStudy (a_clan):

all real except 3, -1

OpenStudy (anonymous):

x cant be 3 or -1

OpenStudy (anonymous):

In domain problems, try to find out which solutions will you give infinity or underterminate answers, and rule them out!

OpenStudy (anonymous):

Wow. Would you mind showing me how you got that answer?

OpenStudy (anonymous):

basicaly this is in form of rational no... p/q...where q must not be 0. so put denominator =0 and find values

OpenStudy (anonymous):

and those values won't be in domain as those give q =0

OpenStudy (a_clan):

Any one of the terms (x-3) or (x+1) becoming zero will make the fraction undefined(division by zero). So, x cannot be 3 or -1.

OpenStudy (anonymous):

division by zero is infinity not undefined ( 0/0 is undefined)

OpenStudy (a_clan):

infinity is also undefined

OpenStudy (anonymous):

So it would be all x not equal to 3, x not equal to -1?

OpenStudy (a_clan):

correct

OpenStudy (anonymous):

Thank you so very much!! GOD bless you for your kindness! While a smart person is here, can you help me with another question? Find the equation in general form of the line (5,-4) (-4,3)

OpenStudy (anonymous):

First find the slope

OpenStudy (anonymous):

try findin slope m= (y2 -y1)/x2 -x1 then put in the eq y = mx +c ....then put any of the coordinates and find c.... note . 2 and 1 are subscript